OpenStudy (anonymous):

Word Problem: Pete found a watch for $4.50, Pete sold it for 35 dollars. What was the percent of the increase in price? Round to the nearest hundredth percent.

OpenStudy (mathstudent55):

\( \bf percent ~change = \dfrac{new ~number - old ~number}{old ~number} \times 100 \) If the change is positive, it is a percent increase (what you have here). If the change is negative, it is a percent decrease.

OpenStudy (anonymous):

could you break it down more in numbers

OpenStudy (mathstudent55):

\( percent ~change = \dfrac{$35 - $4.50}{$4.50} \times 100 \)

OpenStudy (anonymous):

cool, thanks

OpenStudy (mathstudent55):

wlcm

OpenStudy (anonymous):

Jon bought a Ford truck with a down payment of 2,000, he still owes 80% of the selling price, what was the selling price of the truck?

OpenStudy (anonymous):

Help please anyone to the question I posted previously

OpenStudy (anonymous):

Word Problem: Alice made a 3,000 down payment of a new Ford Explorer, she still owes 90% of the selling price. What was the selling price of the wagon?

OpenStudy (mathstudent55):

For this problem: Jon bought a Ford truck with a down payment of 2,000, he still owes 80% of the selling price, what was the selling price of the truck?

OpenStudy (mathstudent55):

If he still owes 80%, then the down payment is 20%, right?

OpenStudy (anonymous):

ok, explain

OpenStudy (mathstudent55):

The total price of the car is an unknown amount. It is 100% of the price of the truck. 100% means the entire thing.

OpenStudy (mathstudent55):

Since the entire price is 100%, and he still owes 80%, then 100% - 80% = 20%. The amount he paid is 20% of the price.

OpenStudy (mathstudent55):

Total price: 100% He owes: 80% ---------------------(subtract) He paid: 20%

OpenStudy (mathstudent55):

Now we know that he paid 20% of the total price. We are told that he paid $2000. That means 20% of the total price is $2000. We use a proportion: 20% is to $2000 as 100% is to x \(\dfrac{20}{2000} = \dfrac{100}{x} \) 20x = 2000 * 100 20x = 200000 x = 10000 The price of the truck is $10,000

OpenStudy (mathstudent55):

For this problem: "Word Problem: Alice made a 3,000 down payment of a new Ford Explorer, she still owes 90% of the selling price. What was the selling price of the wagon?" This is similar to the above problem just with different numbers. If she owes 90%, then 100% - 90% = 10%. That means her down payment was 10%. 10% is to 3000 as 100% is to x. \(\dfrac{10}{3000} = \dfrac{100}{x} \) 10x = 100 * 3000 10x = 300000 x = 30000 Answer: The price is $30,000

OpenStudy (anonymous):

thank you so much you make math fun to learn especially word problems

OpenStudy (anonymous):

Word Problem: In 2010, Dairy Queen had 550,000 in sales. In 2011, Dairy Queen's sales were up 35% What were Dairy Queens sales in 2011

OpenStudy (mathstudent55):

You can do this 2 ways. I'll show you both ways, and in the future you can pick the way you like better. Method 1: Find what 35% of $550,000 is and add to $550,000 35% of $550,000 = 0.35 x $550,000 = $192,500 35% of $550,000 is $192,500 Now we add that to $550,000 $550,000 + $192,500 = $742,500 The answer is $742,500 Method 2: The sales of 2010, were 100% of the sales. Since the sales for 2011 were up 35%, that means in 2011, the sales were 135% of what they were in 2010 (100% + 35% = 135%). 135% of $550,000 = 1.35 x $550,000 = $742,500 The answer is $742,500 As you can see, you get the same final answer using either method.